A thermometer is taken from a room where the temperature is 20°C to the outdoors,

where the temperature is 5°C. After one minute the thermometer read 12°C.Use

Newton’s Law of Cooling to answer the following questions:

a) What will the reading on the thermometer be after one more minute?

b) When will the thermometer read 6°C?


1
Expert's answer
2022-01-24T16:18:36-0500

As per Newton's Law of Cooling, the rate of the cooling is always proportional to the temperature difference between the object and its surroundings.

"\frac{dT}{dt}=-k(T_0-T_s)"

"T_t=T_s+(T_0-T_s)e^{-kt}"

"T_0=20\u00b0C" And "T_s=5\u00b0C"

At t=60, "T_{60}=12\u00b0C"

"12=5+(20-5)e^{-60k}"

"12-5=15e^{-60k}"

"\frac{7}{15}=e^{-60k}"

"k=-\frac{ln(\frac{7}{15})}{60}"

"k=0.0127023342"

a)

After another one minute, "\therefore t=120s"

"T_t=T_s+(T_0-T_s)e^{-kt}"

"T_{120}=5+(15)e^{-0.0127023342*120}"

"=5+0.01611786424"

"=5.01611786424\u00b0C"


b)

When temperature will read 6°C

"T_t=T_s+(T_0-T_s)e^{-kt}"

"6=5+(15)e^{-0.0127023342t}"

"1=(15)e^{-0.0127023342t}"

"e^{-0.0127023342t}=\frac{1}{15}=0.06666666667"

"t=-\frac{ln 0.06666666667}{0.0127023342}"

t=213.19s

=3.55minutes